#4 and 5 Thank you!
Problem 4 & 5: Find the slope of the graph off at the given point. 4. f(x) = 5 – 3x (1, -2) 5. f(x) = V + 2 (9,5)
Answer #1


Ary.4) frol: 5-30 at (1,-2) Slobo IP -3 doc dois 1-3 Ary 52 ad forl= JXC+2 5 (9,5) Slobi IP T o+ dx at 1-9 1 6 Siop= 7 - Ary